OpenStudy (anonymous):

At what temperature will the root-mean-square speed of oxygen molecules have the value of 640 m/s? 1 kilomole of oxygen has a mass of 32kg

OpenStudy (anonymous):

v =sqrt(3RT/m) where m is molar mass (in out case 0.032 kg) R is gas constant = 8.3 hence we have T = v^2*m/(3R) = 640^2*0.032/(3*8.3) = 526.4 K = 253.25 C